ON THE ELECTRONIC SPECTROSCOPY OF CLOSED SHELL CATIONS DERIVED FROM RESONANCE STABILIZED RADICALS: INSIGHTS FROM THEORY AND FRANCK-CONDON ANALYSIS

TYLER P. TROY, SCOTT H. KABLE, TIMOTHY W. SCHMIDT, School of Chemistry, The University of Sydney, NSW 2006, Australia; \underline{SCOTT A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

Recent attention has been directed on closed shell aromatic cations as potential carriers of the diffuse interstellar bands. The spectra of mass-selected, matrix-isolated benzylium and tropylium cations were recently reported [Nagy, A., Fulara, J., Garkusha, I. and Maier, J. P. (2011), Angew. Chem. Int. Ed., 50: 3022-3025]. The benzylium spectrum shows an extended progression in a low frequency (510 cm^{−1}) ring distortion mode. Modeling of the benzylium spectrum using (TD)DFT and MCSCF-PT2 methods in concert with multidimensional Franck-Condon (FC) analysis is found to yield excellent agreement with the experimental spectrum. We extended this analysis to larger (2 and 3 ring) PAH cations derived from resonance stabilized radicals, which are predicted to show strong S_{0} → S_{n} transitions in the visible region. The FC progression is significantly quenched in the larger species, and our results for 1-napthylmethylium are in excellent agreement with very recent experiments [Nagy, A., Fulara, J., and Maier, J. P. (2011), J. Am. Chem. Soc., 133, 19796]. Since carriers of the DIBs should exhibit spectra dominated by a single vibronic transition, our results demonstrate that closed-shell cations may present spectra with the required properties. Furthermore, the calculated ionization energies of a range of CSCs were found to be in the 13-14 eV range, consistent with variations in behaviour of the DIB carriers with respect to various astrophysical environments.